OpenStudy (anonymous):

Which is an equation in point-slope form for the given point and slope? Point (2,-6) Slope -3/4

OpenStudy (jojo4eva):

y+6=-3/4(x-2)

OpenStudy (anonymous):

You could use the equation y = mx + b, where m is the slop, and the y-intercept is b.

OpenStudy (anonymous):

Can you figure it out now? @Kaso

OpenStudy (jojo4eva):

Its point slope form @AnswerMyQuestions not slop intercept form

OpenStudy (anonymous):

Oh, thank you for telling me that.

OpenStudy (anonymous):

ca you help me with more @jojo4eva

OpenStudy (jojo4eva):

no problem. Didnt mean to offend @AnswerMyQuestions juss wanted to make you aware of what was being asked. And yes @Kaso i can just tag me

OpenStudy (anonymous):

Oh, I wasn't offended. :)